为了减少道路上的超速事件数量,S国的交警在三个繁忙的交通路口 A、B 和 C 设置了交通摄像头,以监控通过这些路口的车辆速度。路口 A、B 和 C 的摄像头捕获的超速车辆平均数分别为每 3 小时 2 辆、每 4 小时 5 辆和每小时 λ\lambda 辆。假设这三个路口的摄像头捕获的超速车辆数服从泊松分布。

  1. (a)

    求在一小时内,路口 A 和 B 捕获的超速车辆总数至少为 2 辆的概率。

  2. (b)

    已知在一小时内,这三个路口共捕获了 2 辆超速车辆,求其中至少有一辆超速车辆是在路口 C 被捕获的概率。答案用 λ\lambda 表示。

  3. (c)

    已知交通摄像头全天 24 小时运行,使用适当的近似方法,求在一天内路口 A 捕获的超速车辆比路口 B 多的概率。写出使该计算合理的一个假设。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

9(a)

1. Define the combined random variable for the number of vehicles caught at Junctions A and B in an hour.

Let XX and YY be the number of speeding vehicles caught at Junctions A and B in an hour respectively.
X∼Po(23)X \sim \text{Po}(\frac{2}{3})
Y∼Po(54)Y \sim \text{Po}(\frac{5}{4})
Assuming the number of vehicles caught at each junction are independent, let W=X+YW = X + Y be the total number of vehicles caught at both junctions.
W∼Po(23+54)W \sim \text{Po}(\frac{2}{3} + \frac{5}{4})
W∼Po(2312)W \sim \text{Po}(\frac{23}{12})

2. Calculate the probability of at least 2 speeding vehicles being caught.

P(W≥2)=1−P(W≤1)\text{P}(W \ge 2) = 1 - \text{P}(W \le 1)
P(W≥2)=1−(P(W=0)+P(W=1))\text{P}(W \ge 2) = 1 - \left( \text{P}(W = 0) + \text{P}(W = 1) \right)
P(W≥2)=1−e−2312(1+2312)\text{P}(W \ge 2) = 1 - e^{-\frac{23}{12}} \left( 1 + \frac{23}{12} \right)
P(W≥2)≈1−0.42903≈0.571 (to 3 s.f.)\text{P}(W \ge 2) \approx 1 - 0.42903 \approx 0.571 \text{ (to 3 s.f.)}

答案

0.5710.571

9(b)

1. Define the Poisson parameters for each junction and the combined parameter for Junctions A and B.

Let XAX_A, XBX_B, and XCX_C be the number of speeding vehicles caught in an hour at Junctions A, B, and C respectively.
XA∼Po(23)X_A \sim \text{Po}\left(\frac{2}{3}\right)
XB∼Po(54)X_B \sim \text{Po}\left(\frac{5}{4}\right)
XC∼Po(λ)X_C \sim \text{Po}(\lambda)
Let W=XA+XBW = X_A + X_B be the number of speeding vehicles caught at Junctions A and B in an hour.
Assuming independence, the sum follows a Poisson distribution:
W∼Po(23+54)⇒W∼Po(2312)W \sim \text{Po}\left(\frac{2}{3} + \frac{5}{4}\right) \Rightarrow W \sim \text{Po}\left(\frac{23}{12}\right)

2. Formulate the required conditional probability using the complement rule.

Let Y=W+XCY = W + X_C be the total vehicles caught at all three junctions in an hour, so Y∼Po(2312+λ)Y \sim \text{Po}\left(\frac{23}{12} + \lambda\right).
We need to find P(XC≥1∣Y=2)P(X_C \ge 1 \mid Y = 2).
Using the complement rule:
P(XC≥1∣Y=2)=1−P(XC=0∣Y=2)P(X_C \ge 1 \mid Y = 2) = 1 - P(X_C = 0 \mid Y = 2)
By the definition of conditional probability:
P(XC=0∣Y=2)=P(XC=0∩Y=2)P(Y=2)=P(XC=0∩W=2)P(Y=2)P(X_C = 0 \mid Y = 2) = \frac{P(X_C = 0 \cap Y = 2)}{P(Y = 2)} = \frac{P(X_C = 0 \cap W = 2)}{P(Y = 2)}

3. Calculate the constituent probabilities and simplify the final expression.

Since XCX_C and WW are independent, P(XC=0∩W=2)=P(XC=0)P(W=2)P(X_C = 0 \cap W = 2) = P(X_C = 0)P(W = 2):
P(XC=0)=e−λP(X_C = 0) = \text{e}^{-\lambda}
P(W=2)=e−2312(2312)22!P(W = 2) = \text{e}^{-\frac{23}{12}} \frac{\left(\frac{23}{12}\right)^2}{2!}
P(Y=2)=e−(2312+λ)(2312+λ)22!P(Y = 2) = \text{e}^{-\left(\frac{23}{12} + \lambda\right)} \frac{\left(\frac{23}{12} + \lambda\right)^2}{2!}
Substituting these into the conditional probability ratio:
P(XC=0∣Y=2)=e−λ⋅e−2312(2312)22e−(2312+λ)(2312+λ)22=(2312)2(2312+λ)2=(231223+12λ12)2=(2312λ+23)2\begin{aligned} P(X_C = 0 \mid Y = 2) &= \frac{\text{e}^{-\lambda} \cdot \text{e}^{-\frac{23}{12}} \frac{\left(\frac{23}{12}\right)^2}{2}}{\text{e}^{-\left(\frac{23}{12} + \lambda\right)} \frac{\left(\frac{23}{12} + \lambda\right)^2}{2}} \\ &= \frac{\left(\frac{23}{12}\right)^2}{\left(\frac{23}{12} + \lambda\right)^2} \\ &= \left( \frac{\frac{23}{12}}{\frac{23 + 12\lambda}{12}} \right)^2 \\ &= \left( \frac{23}{12\lambda + 23} \right)^2 \end{aligned}
Therefore, the required probability is:
P(XC≥1∣Y=2)=1−(2312λ+23)2P(X_C \ge 1 \mid Y = 2) = 1 - \left( \frac{23}{12\lambda + 23} \right)^2

答案

1−(2312λ+23)21 - \left( \frac{23}{12\lambda + 23} \right)^2

9(c)

1. Determine the distribution of the number of speeding vehicles caught at Junctions A and B in a day.

Let AA and BB be the number of speeding vehicles caught in a 24-hour day at Junction A and Junction B respectively.
For Junction A: λA=24×23=16\lambda_A = 24 \times \frac{2}{3} = 16
For Junction B: λB=24×54=30\lambda_B = 24 \times \frac{5}{4} = 30
Thus, A∼Po(16)A \sim \text{Po}(16) and B∼Po(30)B \sim \text{Po}(30).

2. State the normal approximations for $A$ and $B$.

Since λA=16>10\lambda_A = 16 > 10 and λB=30>10\lambda_B = 30 > 10, the Poisson distributions can be approximated by Normal distributions:
A∼N(16,16)A \sim \text{N}(16, 16) approximately.
B∼N(30,30)B \sim \text{N}(30, 30) approximately.

3. Find the distribution of the difference $A - B$.

Let W=A−BW = A - B.
Assuming AA and BB are independent,
E(W)=E(A)−E(B)=16−30=−14\text{E}(W) = \text{E}(A) - \text{E}(B) = 16 - 30 = -14
Var(W)=Var(A)+Var(B)=16+30=46\text{Var}(W) = \text{Var}(A) + \text{Var}(B) = 16 + 30 = 46
Therefore, W∼N(−14,46)W \sim \text{N}(-14, 46) approximately.

4. Apply a continuity correction to find the required probability.

We require P(A>B)\text{P}(A > B), which is equivalent to P(A−B>0)\text{P}(A - B > 0).
Since A−BA - B takes integer values, P(A−B>0)=P(A−B≥1)\text{P}(A - B > 0) = \text{P}(A - B \ge 1).
Applying the continuity correction for the normal approximation, we calculate P(W>0.5)\text{P}(W > 0.5).
Using a graphing calculator:
P(W>0.5)≈0.016263≈0.0163\text{P}(W > 0.5) \approx 0.016263 \approx 0.0163 (to 3 s.f.)

5. State the necessary assumption for the calculation to be valid.

Assumption: The number of speeding vehicles caught at Junction A and the number caught at Junction B are independent of each other (this is required to sum their variances).

答案

0.01630.0163
Assumption: The number of speeding vehicles caught at Junction A is independent of the number of speeding vehicles caught at Junction B.

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5已知事件 AA 和 BB 满足 P(A)=0.7P(A)=0.7 且 P(B)=0.8P(B)=0.8。证明 P(A∩B)≥0.5P(A \cap B) \ge 0.5 若 P(A′ ∪ B)=0.85P(A'\ \cup\ B)=0.85,求 P(A∣B)P(A\mid B)。说明并解释 AA 和…